OpenStudy (anonymous):

Simplify. sqrt(x) = square root 3/sqrt(2) - 5i

OpenStudy (anonymous):

Oh and i = imaginary number

geerky42 (geerky42):

\[\dfrac{3}{\sqrt 2} - 5i\] ?

OpenStudy (anonymous):

Nope, -5i is also in the denominator

OpenStudy (anonymous):

\[\frac{ 3 }{ \sqrt{2} -5i }\]

OpenStudy (acxbox22):

start off by rationalizing the denominator

OpenStudy (anonymous):

Simple English please :D

geerky42 (geerky42):

Ah okay. \[\dfrac{3}{\sqrt 2 - 5i}\] So first thing to do is multiply both numerator and denominator by conjugate of denominator (\(\sqrt2 + 5i\)) Like this: \[\dfrac{3}{\sqrt 2 - 5i}\cdot\dfrac{\sqrt2 + 5i}{\sqrt2 + 5i}\]

OpenStudy (anonymous):

√2+5ı/9

geerky42 (geerky42):

\[\cdots=\dfrac{3(\sqrt2 + 5i)}{(\sqrt{2})^2-(5i)^2}\]

geerky42 (geerky42):

Now you just have integer in denominator. That's how you simplify.

geerky42 (geerky42):

Does that make sense? @voce

OpenStudy (anonymous):

I got this as the final answer\[\frac{ 3\sqrt{2}+15i }{ 17 }\]

geerky42 (geerky42):

You mean \[\frac{ 3\sqrt{2}+15i }{ 27 }\] You still have to simplify more. Note that numerator and denominator is divisible by 3.

geerky42 (geerky42):

\[\frac{ 3(\sqrt{2}+5i) }{ 3\cdot9 }\]

OpenStudy (anonymous):

OH okay got you, thanks for your help

geerky42 (geerky42):

You are welcome